OpenStudy (anonymous):

What is the x value for the system of equations? 2x + 3y = 4 x - 2y = -5 ---- I got y = 2 but when i put that into the first equation I ended up with x = 0. However, my answer choices are -1, 1, 2, and -2

OpenStudy (harsimran_hs4):

you can use substitution method 1. find the value of y in terms of x from 2nd equation and tell what you get ?

OpenStudy (johnweldon1993):

We want the 'x' value? okay....so lets solve that top equation for 'y' 2x + 3y = 4 subtract 2x from both sides 3y = -2x + 4 divide everything by 3 y = -2/3x + 4/3 Now substitute this in for 'y' in that second equation x - 2y = -5 Will become x - 2(-2/3x + 4/3) = -5 distribute the -2 x + 4/3x - 8/3 = -5 Add 8/3 to both sides x + 4/3x = -5 + 8/3 well x + 4/3x = 7/3x and -5 + 8/3 = -7/3 So we have 7/3x = -7/3 x = -1 is what I get *elimination method here would have been easier*

OpenStudy (anonymous):

2x + 3y = 4 -->(2) x - 2y = -5 --> (3) ------------ 4x + 6y = 8 (result of multiplying by 2) 3x - 6y = -15 (result of multiplying by 3) ------------add 7x = -7 x = -1